# Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents

### v1

@@ 3.2. Experiment Setup
We devised the following experimental scenario to investigate human cooperative behavior patterns when interacting with LLM agents functioning as potential independent decision-makers in human society. The experiment consisted of four sessions. Before each session, participants were informed about the characteristics of their opponents (you can also think of it as opponents’ nature), which were, respectively, a purported human, a purported rule-based AI agent, an LLM agent, and another participant. However, the actual situations are in the Table  1.
Before the commencement of the experiment, participants received an introduction to the characteristics of the different opponents and scoring rules. This introduction was designed to ensure that participants fully understood the nature of the rule-based AI agents and LLM agents, as highlighted in Silva et al. (2023); Igwe and Durrheim (2024). All experiments ensured that participants comprehended the types of AI agents employed in the study. Ensuring this understanding is crucial for the validity of the behavior experiment, as it allows participants to interact appropriately with the various AI agents, thereby guaranteeing the accuracy and reliability of our experimental results.
The core description of the rule-based AI agent highlighted that is based on human-coded rules and algorithms to guide their behavior in gameplay. Expert programmers carefully design these algorithms to ensure the agent’s decision-making process is interpretable, predictable, and grounded in logical reasoning (Russell and Norvig, 2016). However, we cannot disclose the underlying logic and code structure to participants. For the agents driven by LLMs, we emphasized that these models were neither pre-trained for the experiment nor subjected to human manipulation. Additionally, we informed participants that their LLM agent opponents are based on the ChatGPT API [^3] https://www.openai.com, specifically using the gpt-4-turbo-20240409 version. This emphasis was intended to assure participants that no biases or prior knowledge bases were imposed to enhance the agents’ understanding of the game. Bender and Koller (2020) highlights that fine-tuning is often specific to particular experimental contexts and can easily result in model overfitting. Given that our research emphasizes human behavior, utilizing an unfine-tuned model is more suitable to ensure the reliability and generalizability of our findings. Furthermore, it aims to replicate real-world interactive environments, where pre-training LLM agents for every possible scenario is impractical (Schick and Schütze, 2020).
Participants were arranged in standardized scenarios during the experiment, depicted in Figure 1. They were given two cards on the table, labeled “Cooperation” and “Defection”. They were asked to use their fingers to indicate their choices. During the first three trials, the room designated for the other associate was unoccupied, as the LLM agent was operating on the researcher’s laptop. Each participant arrived at the experimental site separately, ensuring their experiences were isolated. Upon receiving the choices—either cooperation or defection—from both the human participant and the LLM agent for each round, the researcher communicated the choices to both parties (dictating them to the participant and inputting the results for the LLM agent). After the first three trials, all participants completed a five-question questionnaire and participated in a 15-minute semi-structured interview one by one.
Within two days of each participant completing the initial three trials, they were invited again to complete the final trial. This last trial was distinctive because it involved two participants who had previously participated in the initial three rounds. These participants were invited to separate rooms and given strict instructions prohibiting communication. To ensure the privacy of each participant, they were directed to the experimental rooms via different corridors, ensuring they would not encounter each other upon leaving. Figure 2 indicates the experiment process.
@@ 3 LLM Agent LLM Agent 50
4 Another Participant Another Participant 50
3.3. Polit Experiment
To ensure our experimental design is accessible, we conducted a pilot study involving 8 participants (four males and four females) before the formal experiment. Initially, each participant was required to complete four consecutive game trials in a single session (purported human, purported rule-based AI agents, LLM agents, and another participant). This setup, however, resulted in noticeable fatigue and boredom among the participants during the experiment, which caused them to abandon strategy variation and persist with the same choices throughout the trials. Furthermore, we observed that participants were significantly more cooperative with another human participant than with other agents. We carefully considered the possibilities, and the elevated cooperation rate was likely driven by the monetary reward, which strongly incentivized collaborative behavior. Furthermore, a preliminary data analysis from these 8 participants revealed a potential influence of gender and the AI agent’s purported characteristics on cooperative behavior. Building on these insights, we designed a formal experiment and recruited a new cohort of participants.
In response to these findings, we made several critical adjustments to the main experiment. We divided the experimental sessions into two separate phases. Participants interacted with three AI agents in one phase, and in the other, they played with another human participant. Specifically, participants were provided with blank sheets of paper to record these choices during the games, facilitating accurate recall during subsequent questionnaires and interviews. These modifications aimed to mitigate participant fatigue, enhance engagement, and improve the overall validity and reliability of the experimental outcomes.
3.4. Game Rules
All participants (N = 30) completed the prisoner’s dilemma game four times across two separate in-person sessions within two days; each consisted of 50 successive rounds, totaling 200 rounds. Based on the original rules of the prisoner’s dilemma, we adapted the model to incorporate a rewarding mechanism, as illustrated in Table 2. Participants were expected to maximize their scores during the game.
@@ 3.4. Game Rules
This adaptation retains the same payoff structure for all four times of the game. The total score $S$ of 200 rounds for each participant was calculated and multiplied by a fixed coefficient $k$ to arrive at the compensation cost $C$. Specifically, for each participant $i$, the compensation cost $C_{i}$ was calculated as follows:
@@
This method ensured that all participants could receive real cash compensation based on their actions, thereby motivating them to engage earnestly with the game. Furthermore, this approach more accurately simulates the incentives and benefits associated with real-world scenarios, thereby enhancing the ecological validity of the experiment. To ensure that participants fully comprehended the scoring rules, detailed instructions were provided for them to read before the commencement of the experiment. To verify their understanding, the trial began with two hypothetical scenarios, and participants were asked to calculate the points awarded in each situation. The experiment officially commenced after all participants correctly answered these preliminary questions.
3.5. LLM Agents Settings
@@
4. User study
4.1. Participants
Participants with experience using LLM-supported AI agents (e.g., code assistants and chat boxes) were recruited for the experiment. Eventually, we have thirty (15 males, 15 females) eligible participants with demographic information presented in Table 3, with ages ranging from 18 to 29, $Mean_{age}$ = 23.7, $SD_{age}$ = 2.38. Regarding the usage of LLM applications, 23 participants (76.7%) reported using LLM applications almost every day, 6 participants (20.0%) used them weekly, and 1 participant (3.3%) used them monthly. The most frequently mentioned LLM applications were ChatGPT and Copilot Code, indicating a preference for these tools among the sample. Before recruitment, we obtained approval from the Ethics Department of xxxxxxxx University and the Ethics Compliance Review, protocol No. xxxxxxxxxx.
Table 3. An Overview of Participants’ Demographics and Freq. of LLMs Usage
ID Gender Age Freq. ID Gender Age Freq.
M1 Male 23 Daily M2 Male 23 Daily
M3 Male 23 Daily M4 Male 22 Weekly
M5 Male 24 Daily M6 Male 26 Weekly
M7 Male 23 Daily M8 Male 22 Daily
M9 Male 29 Daily M10 Male 24 Daily
M11 Male 18 Daily M12 Male 23 Daily
M13 Male 23 Weekly M14 Male 23 Daily
M15 Male 25 Daily F1 Female 24 Daily
F2 Female 22 Daily F3 Female 25 Daily
F4 Female 23 Weekly F5 Female 24 Monthly
F6 Female 25 Daily F7 Female 24 Daily
F8 Female 28 Weekly F9 Female 23 Daily
F10 Female 22 Daily F11 Female 24 Weekly
F12 Female 21 Daily F13 Female 24 Daily
F14 Female 29 Daily F15 Female 23 Daily
@@ 4.2. Measurement
Within the framework of our experimental design, two independent variables were meticulously delineated: the participants’ gender (classified as male or female) and the purported characteristics of the opponents, which encompassed purported human opponents, purported rule-based AI agents, and LLM agents. We defined the following metric to be analyzed from the user study:
@@
  • Cooperation Rate, which refers to the proportion of participants’ choices to cooperate when confronted with opponents of varying purported characteristics in each set of 50 consecutive rounds of the game.
@@
  • Contemplation Time, which denotes the duration participants took to deliberate decisions under identical game conditions in each set of 50 consecutive rounds of the game.
  • Amicability, which pertains to the extent to which humans are willing to initiate cooperation and demonstrate friendliness, either before establishing bilateral cooperation or after a breach in cooperation. It is quantified by the number of rounds in which the participant chooses to cooperate with their opponent from the point of mutual cooperation (or from the last instance of such cooperation if there is a consecutive sequence) until one party breaks this cooperation, and then up until the next occurrence of mutual cooperation. Calculations commence from the tenth round because participants are typically in an exploratory phase during the initial rounds. This approach facilitates strategy stabilization and reduces noise in the data. Early rounds typically involve participants learning the game mechanics and experimenting with different strategies (Axelrod and Hamilton, 1981), which can result in high variability and unstable decision-making. Nowak and Sigmund (1993) also emphasizes the importance of eliminating exploratory behavior to capture genuine strategic behavior. As we demonstrate in Section 2.2, this metric is particularly relevant because, in the period following the breakdown of cooperation and leading up to its reestablishment, if a participant chooses to cooperate, it is considered a gesture of amicability to reinitiate a collaborative relationship. In this context, the figure we refer to is obtained by summing all attempts at amicability and then dividing by the total number of rounds that elapsed between these attempts.
@@
  • Repair Effort, which addresses the degree of conciliatory gesture required from an opponent to obtain human forgiveness and reinstate a cooperative relationship after a breakdown in cooperation, or to demonstrate friendliness before any bilateral cooperation has been established, which can also be understood as the opponent’s “Amicability”. It is calculated in the same manner as “Amicability”, focusing on the number of cooperative actions initiated by the participants’ opponents. This variable aims to measure the extent to which human players accept gestures of friendliness from opponents with different purported characteristics.
@@ 5. Results
We employed a Linear Mixed Effects (LME) model for data analysis (Kuznetsova et al., 2017). This approach enabled us to assess significant effects by examining fixed effects, including the primary effects of participants’ gender and the purported characteristics of the opponents, as well as their interaction (Rosnow and Rosenthal, 1989). These fixed effects were scrutinized to understand their influence on participants’ cooperative behavior when interacting with AI agents in competitive environments.
In addition to the fixed effects, we incorporated participant user ID as a random intercept to account for inherent individual differences that could independently affect the outcomes beyond the fixed effects (Barr, 2013). Initially, we attempted to include random slopes for opponent characteristics to capture individual variability in responses to experimental manipulations across different opponents. However, the multilevel structure of the opponent characteristics resulted in an overparameterized model that failed to converge due to the limited number of observations per participant (Bates et al., 2015). Consequently, we opted for a simpler random intercept model, effectively accounting for individual differences while ensuring model reliability and interpretability.
Table 4. Effects of Gender and Purported Characteristics
Effect Cooperation Time Amicability Repair
@@
Gender .2195 .6428 - 13.3980 <.001 *** .4201 .5218 - .9216 .3447 -
Purported Characteristic 4.5979 .0139 * 30.8168 <.001 *** 1.4950 .2324 - 4.2045 .0196 *
Gender:Purported Characteristic 9.2411 <.001 *** 4.2456 .0189 * 5.5060 .0064 ** .4749 .6243 -
In the table above, asterisks (*) are used to denote the level of statistical significance of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
@@
5.1. Behind the Masks: Cooperation Behavior
Table 5. Characteristic Comparisons: Significance and Marginal Means
Comparison Estimate t Ratio Pr(>F) Sig
@@ 5.1. Behind the Masks: Cooperation Behavior
Purported Human vs. LLM Agents .0633 2.234 .0291 *
Purported Rule-Based AI Agents vs. LLM Agents -.0013 -.047 .9626 -
Purported Characteristic Emmean Lower CL Upper CL
Purported Human .540 .489 .591
Purported Rule-Based Agents .475 .425 .526
Purported LLM Agents .477 .426 .527
In the table above, asterisks (*) denote the statistical significance level of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
Figure 3. Cooperation Rate by Purported Opponents’ Characteristic
The selection of the 4th-degree polynomial model is supported by its superior performance metrics, including the highest $R^{2}$ (0.677) and $\text{Adj}_{R^{2}}$ (0.668), as well as the lowest AIC (-288.122), BIC (-270.058), and $\text{CV}_{\text{RMSE}}$ (0.090) (Vrieze, 2012).
We constructed another linear mixed-effects model using only the purported characteristic as the predictor, and the results are presented in Table 5. Participants exhibited a higher cooperation rate with agents characterized as purported humans, showing a significant difference compared to rule-based AI agents ($p$ = .0260) and LLM agents ($p$ = .0291). However, no significant difference ($p$ = .9626) was observed between rule-based AI agents and LLM agents. Figure 3 further visualizes the variation in cooperation rates across different rounds, highlighting significant disparities.
To examine interaction effects, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics. Pairwise comparisons of these EMMs identified differences in cooperative behavior across genders and opponent characteristics. This method enabled us to assess the significance of differences between each pair and obtain more nuanced insights into detailed results.
Figure 4. Completion Time by Purported Opponents’ Characteristic
Table 6. Interaction Effect Results on Cooperation

### v2

@@ 3.1.1. Declared Identity Setting
We devised the following experimental scenario to investigate human cooperative behavior patterns when interacting with LLM agents functioning as potential independent decision-makers in human society. The experiment consisted of three sessions. Before each session, participants were informed about the identities of their opponents, which were, respectively, a purported human, a purported rule-based AI agent, an LLM agent. However, the actual situations are in the Table  1.
@@ 2 LLM Agent LLM Agent 50
3 Real Human LLM Agent 50
For situation 1, the core description of the rule-based AI agent highlighted that is based on human-coded rules and algorithms to guide their behavior in gameplay. Expert programmers carefully design these algorithms to ensure the agent’s decision-making process is interpretable, predictable, and grounded in logical reasoning (Russell and Norvig, 2016). However, we cannot disclose the underlying logic and code structure to participants. For situation 2, the agents driven by LLMs were emphasized that these models were neither pre-trained for the experiment nor subjected to human manipulation. Additionally, we informed participants that their LLM agent opponents are based on the ChatGPT API [^2] https://www.openai.com, specifically using the gpt-4-turbo-20240409 version. This emphasis was intended to assure participants that no biases or prior knowledge bases were imposed to enhance the agents’ understanding of the game. Besides, Bender and Koller (2020) highlights that fine-tuning is often specific to particular experimental contexts and can easily result in model overfitting. Given that our research emphasizes human behavior, utilizing an unfine-tuned model is more suitable to ensure the reliability and generalizability of our findings. Furthermore, it aims to replicate real-world interactive environments, where pre-training LLM agents for every possible scenario is impractical (Schick and Schütze, 2020).
For situation 3, as a baseline setting for comparing humans’ attitude towards LLM agents, participants were explicitly informed that their opponent represented a real human player in this experimental condition. No demographic particulars (including gender, age, or social identity) were disclosed to preserve anonymity and prevent potential interpersonal bias. The human experimenter assigned to simulate this role maintained complete anonymity throughout the study - no physical, verbal, or digital interaction occurred between participants and the role-playing researcher at any experimental phase (pre-test, during, or post-test).
3.1.2. LLM Agents Settings
3.1.3. Game Rules
All participants should complete the prisoner’s dilemma game three times in in-person sessions based on a random sequence, with LLM agents with different purported roles stated in Table 1; each consisted of 50 successive rounds, totaling 150 rounds. Based on the original rules of the prisoner’s dilemma, we adapted the model to incorporate a rewarding mechanism, as illustrated in Table 2. Participants were expected to maximize their scores during the game.
@@ 3.1.3. Game Rules
This adaptation retains the same payoff structure for all four times of the game. The total score $S$ of 150 rounds for each participant was calculated and multiplied by a fixed coefficient $k$ to arrive at the compensation cost $C$. Specifically, for each participant $i$, the compensation cost $C_{i}$ was calculated as follows:
@@
This method ensured that all participants could receive real cash compensation based on their actions, thereby motivating them to engage earnestly with the game. Furthermore, this approach more accurately simulates the incentives and benefits associated with real-world scenarios, thereby enhancing the ecological validity of the experiment.
To ensure that participants fully comprehended the scoring rules, detailed instructions were provided for them to read before the commencement of the experiment. To verify their understanding, the trial began with two hypothetical scenarios, and participants were asked to calculate the points awarded in each situation. The experiment officially commenced after all participants correctly answered these preliminary questions.
3.1.4. Pilot Study
To ensure the accessibility and robustness of the experimental design, a pilot study was conducted with 8 participants (4 males and 4 females) prior to the formal experiment. Preliminary data analysis from the pilot revealed a potential influence of gender and the AI agents’ purported characteristics on cooperative behavior, which ensured the distinct feature of the following experiment findings to some extent. Based on these findings, the formal experiment was refined, and a new cohort of participants was recruited. To enhance accuracy, participants were provided with blank sheets of paper to record their choices during the games, ensuring reliable recall during subsequent questionnaires and interviews. These adjustments were intended to improve participant engagement while bolstering the validity and reliability of the experimental outcomes.
3.2. Formal Experiment
3.2.1. Participants
Participants with experience using LLM-supported AI agents (e.g., code assistants and chat boxes) were recruited for the experiment. Eventually, we have thirty (15 males, 15 females) eligible participants with ages ranging from 18 to 29, $Mean_{age}$ = 23.7, $SD_{age}$ = 2.38. Regarding the usage of LLM applications, 23 participants (76.7%) reported using LLM applications almost every day, 6 participants (20.0%) used them weekly, and 1 participant (3.3%) used them monthly. The most frequently mentioned LLM applications were ChatGPT and Copilot Code, indicating a preference for these tools among the sample. Before recruitment, we obtained approval from the Ethics Department of the local University and the Ethics Compliance Review, protocol No. HKUST(GZ)-HSP-2024-0021.
3.2.2. Experimental Process
Figure 1 indicates the experiment process. Before the commencement of the experiment, participants received an introduction to the characteristics of the different opponents and scoring rules. This introduction was designed to ensure that participants fully understood the nature of the rule-based AI agents and LLM agents, as highlighted in Silva et al. (2023); Igwe and Durrheim (2024). All experiments ensured that participants comprehended the types of AI agents employed in the study. Ensuring this understanding is crucial for the validity of the behavior experiment, as it allows participants to interact appropriately with the various AI agents, thereby guaranteeing the accuracy and reliability of our experimental results.
@@ 3.2.2. Experimental Process
Participants and a experimenter were arranged in standardized scenarios during the experiment, depicted in Figure 2. They were given two cards on the table, labeled “Cooperation” and “Defection” and strict instructions prohibiting communication. They were asked to use their fingers to indicate their choices or the LLM agent’s choices. During the three trials, the room designated for the other associate was occupied with the LLM agent operating on the experimenter’s laptop, while the experimenter sat in front of it for compressing the LLM agent’s decision, inputting the results to the LLM agent and playing the purported role of a real human on the third trial. Participants were situated in the experimental room separately for each trial, ensuring their experiences were isolated and unknown to the LLM agent’s true identity in the other experimental room. Upon receiving the choices—either cooperation or defection—from both the human participant and the LLM agent for each round, another experimenter communicated the choices to both parties (dictating the results to the participant and the LLM agent). After the three trials, participants completed a five-question questionnaire and participated in a 15-minute semi-structured interview.
Figure 2. Experimental Scenarios. (1) The overall layout of the experimental site consists of two glass rooms, where participants can face the experimenter directly; (2) One of the separate rooms; (3) The selection cards for the participants and the provided record sheets.
@@ 3.3. Measurement
3.3.1. Quantitative Metrics
We defined the following metric to describe the cooperation attitude and behaviors during the game process quantitatively:
@@ 3.3.1. Quantitative Metrics
  • Cooperation Rate, which refers to the proportion of participants’ decisions to cooperate when interacting with the opponents, measured across each set of 50 consecutive rounds of the game.
@@
  • Decision Latency, which represents the amount of time participants spent deliberating their decisions under identical game conditions, measured across each set of 50 consecutive rounds of the game.
  • Unsolicited Cooperation Acts, which measures the willingness of participants to initiate cooperation or display friendliness, either before mutual cooperation is established or after a breach in cooperation, as discussed in Section 2.1. This metric is calculated by counting the number of rounds in which a participant chooses to cooperate with their opponent, starting from the last point of mutual cooperation until one party breaks the cooperation (or the last instance of mutual cooperation within a sequence), and continuing up to the next occurrence of mutual cooperation.
@@
  • Trust Restoration Tolerance, which addresses the degree of conciliatory gesture required from an opponent to obtain human forgiveness and reinstate a cooperative relationship after a breakdown in cooperation, or to demonstrate friendliness before any bilateral cooperation has been established, as discussed in Section 2.1. It focuses on the number of cooperative actions initiated by the participants’ opponents. This variable aims to measure the extent to which human players accept gestures of friendliness from opponents with different declared identities.
Notably, to remove noise and ensure strategy stabilization, the calculations of Unsolicited Cooperation Acts and Trust Restoration Tolerance begin from the tenth round, as participants are typically in an exploratory phase during the initial rounds. Early rounds often involve learning game mechanics and experimenting with strategies (Axelrod and Hamilton, 1981), leading to high variability and unstable decision-making. Nowak and Sigmund (1993) similarly emphasizes the importance of excluding exploratory behavior to capture genuine strategic interactions. Their final value is determined by summing all such unsolicited cooperation acts or trust restoration tolerance and dividing by the total number of rounds that elapsed between these acts.
3.3.2. Questionnaires
This questionnaire was designed to probe specific aspects of attitudes and behavior, particularly in situations involving trust, collaboration, and moral judgment, which consists of 5 multiple choice questions with two options “Purported Rule-based AI Agent” and “LLM Agent”. They assist the study in better analyzing and distinguishing the perceptions of rule-based AI agents and LLM agents in the context of the prisoner’s dilemma, as well as how these perceptions influence participants’ cooperative attitudes and behaviors.
  • RQ 1: During the game, whom did you ’punish’ or ’reward’ more often?
  • RQ 2: Based on your experience, whom do you find more reliable?
  • RQ 3: Based on your experience, whom do you prefer to work with?
  • RQ 4: During the recent game, whom did you aim to outperform in terms of score?
  • RQ 5: When someone betrays you, whom do you subject to more moral condemnation?
3.3.3. Semi-structured Interview
We posed two key questions for the interview:
  • What was your strategy against the three opponents, and why?
  • How did you feel and adjust your strategy when facing non-cooperative behaviors (defection)?
The key questions focus on participants’ strategic decisions when interacting with AI agents of varying declared identity and their responses to non-cooperative behaviors. Participants reviewed their recorded game data to aid recollection and contextualize their strategic choices. All interviews were audio-recorded in English, and lasted 10 minutes per participant.
3.4. Data Analysis
We utilized a Linear Mixed Effects Model (LMEM) (Kuznetsova et al., 2017) considering the repetitiveness of the data measurement. This method allowed us to investigate significant effects by examining the fixed effects, including the declared identity of the AI agents, participant’s gender, and their interaction (Rosnow and Rosenthal, 1989). In addition to the fixed effects, we incorporated participant user ID as a random intercept to account for inherent individual differences that could independently affect the outcomes beyond the fixed effects (Barr, 2013). Initially, we attempted to include random slopes for opponent characteristics to capture individual variability in responses to experimental manipulations across different opponents. However, the multilevel structure of the AI agent’s declared identity resulted in an overparameterized model that failed to converge due to the limited number of observations per participant (Bates et al., 2015). Consequently, we opted for a simpler random intercept model, effectively accounting for individual differences while ensuring model reliability and interpretability.
Besides, semi-structured interviews were transcribed. During the transcription process, nonessential fillers and disfluencies were removed. The transcribed data were manually coded using spreadsheets, highlighting tools, and affinity mapping, employing a thematic coding process. To minimize bias, two experimenters collaboratively conducted two rounds of initial coding. The coding process followed an iterative approach, involving two rounds of coding to merge, split, or refine preliminary codes. After finalizing the codes, a separate meeting was held among three experimenter to consolidate these basic codes into themes that directly addressed the research questions. Ultimately, two major themes were identified, encompassing three distinct sub-themes to support the study’s objectives.
@@ 4. Results
We summarized the total results and stated them separately to answer the research questions. As presented in Table 3, we analyzed the effect of declared identity and gender on four metrics, respectively. Furthermore, to further discover the influencing factors within these two variables, we also explored the interactive effects of declared identity and gender.
For all the tables and figures provided in this section: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level.
Table 3. Effects of Gender and AI Agent Declared Identity
Effect Cooperation Rate Decision Latency Unsolicited Cooperation Trust Restoration
@@
Declared Identity 3.3979 .0402 * 26.0963 <.001 *** 1.2211 .3024 - 4.001 .0236 *
Gender .2049 .6542 - 12.5048 .0014 ** .3921 .5362 - .8602 .3616 -
Declared Identity : Gender 8.6250 <.001 *** 3.9626 .0245 * 5.1389 .0089 ** .4432 .6442 -
@@
4.1. The Impact of AI agents’ Declared Identity and Gender
4.1.1. Cooperation Rate: Purported Human > Purported LLM Agents / Purported Rule-based AI Agents (RQ1)
We conducted a linear mixed-effects model to examine the influence of the AI agent’s declared identity on the Cooperation Rate. In this model, the AI agent’s declared identity was included as the sole predictor to isolate its effect on participants’ cooperation rate. As summarized in Table 4, the results reveal that participants exhibited a significantly higher cooperation rate when interacting with agents identified as purported humans, compared to both rule-based AI agents ($\beta$= .0647, t=2.281, $p$ = .0260) and LLM agents ($\beta$= .0633, t=2.234, $p$ = .0291).
Table 4. Declared Identity Contrast on Cooperation Rate
Contrast Estimate t Ratio Pr(>F) Sig
@@ 4.1.1. Cooperation Rate: Purported Human > Purported LLM Agents / Purported Rule-based AI Agents (RQ1)
Purported Human vs. Purported LLM Agents .0633 2.234 .0291 *
Purported Rule-based AI Agents vs. Purported LLM Agents -.0013 -.0470 .9626 -
Purported Human: Emmean = .540, 95%CI = [ .489, .591]; Purported Rule-based Agents: Emmean = .475, 95%CI = [ .425, .526]; Purported LLM Agents: Emmean = .477, 95% = [ .426, .527]
4.1.2. Decision Latency: Purported Human > Purported LLM Agents > Purported Rule-based AI Agents (RQ1)
[…]
Table 5. Declared Identity Contrast on Decision Latency